Solve the equation.
6x-2-7(x+1)=-4x-2

Answer by stanbon(75887)

6x-2-7(x+1)=-4x-2
---
6x-2-7x-7 = -4x-2
---
-x - 9 = -4x-2
3x = 7
x = 7/3
==============
Cheers,
Stan H.
==============